OpenStudy (anonymous):

same problem : solve the following system of equations 5x-2y=1 10x-4y=7

OpenStudy (anonymous):

No solution

OpenStudy (saifoo.khan):

Moneybird copied my answer.

OpenStudy (anonymous):

How? I just logged on

OpenStudy (saifoo.khan):

that's cheating.

OpenStudy (anonymous):

first solve for either x or y in one of your equations, which one do you want to start with?

OpenStudy (anonymous):

i guess x

OpenStudy (anonymous):

ok, so lets look at your first equation, 5x-2y = 1 if we solve this for x that means we need to x on one side by itself, so we need to move the -2y to the other side, which will make it a +2y giving us: 5x = 2y +1

OpenStudy (anonymous):

so now we still don't have x by itself, so we need to divide everything by 5: x = 2/5 y +1/5

OpenStudy (anonymous):

now we know what x is equal to and we can put that into our second equation to get the answer for y. so since x= 2/5 y +1/5, we put that into our second equation as a replacement for x: 10(2/5 y +1/5)-4y=7

OpenStudy (anonymous):

now we can divide 5 into 10 so we can get rid of those gross fractions. that will leave us with: 2(2y+1) - 4y = 7

OpenStudy (anonymous):

now we want to multiply the 2 that was leftover from the 10 into the parenthesis: 2y + 2 -4y = 7

OpenStudy (anonymous):

now we combine like terms and get: -2y + 2 = 7 and solving for y we get: -2y = 5 y= -(5/2)

OpenStudy (anonymous):

now that we know what y is, we can go waaaaaayyyy back to when we figured out what x was equal to and plug it in. so we get: x = 2/5 y +1/5 = (2/5)(-5/2) +(1/5)

OpenStudy (anonymous):

multiplying that out gives us: x= -1 + (1/5)

OpenStudy (anonymous):

which can be simplified as x = -(4/5)

OpenStudy (anonymous):

can you please draw it out its confusing reading it out this way

OpenStudy (anonymous):

so now you have your x and your y and your problem is almost complete! the last thing left is to write them out all nice and neat so that your teacher doesnt miss anything. y= -(5/2) x= -(4/5)

OpenStudy (anonymous):

............... yup.

OpenStudy (anonymous):

thank youuuu

OpenStudy (anonymous):